20. Assume the following information: Exchange rate of Japanese yen in U.S. S Exchange rate of euro in U.S. $ Exchange rate of euro in Japanese yen $.011 $1.40 140 yen What will be the yield for an investor who has $1,000,000 available to conduct triangular arbitrage?
Answer #1

Start with $1,000,000

Convert to Euro at Euro 1.4 dollars per Euro

=1000000/1.4= Euro 714,285.71

Convert o Yen at 140 yen per Euro= 714,285.71*140=Yen 100,000,000

Convert back to USD at 0.011 USD per Yen = 100,000,000*0.011=$1,100,000

Hence profit = $1,100,000-$1,000,000=$100,000

Percentage = $100,000/$1,000,000=10%

Hence yield =10%